# Comparison of f(R,T) Gravity with Type Ia Supernovae Data

### v1

@@ Abstract
The expansion of the universe in $f(R,T)$ gravity is studied. By focusing on functions of the form $f(R,T)=f_{1}(R)+f_{2}(T)$, we assert that present day acceleration can be achieved if the functional form of $f_{2}(T)$ either grows slowly or falls as a function of $T$. In particular, we demonstrate that when $f_{2}(T)\propto T^{-1}$, the universe transitions to exponential growth at late times, just as it does in the standard cosmological model. A comparison of predictions of this model, with Type Ia supernovae shows that this model fits the data as well or even slightly better than the standard cosmological model without increasing the number of parameters.
@@ I Introduction
For over two decades, we have known that the expansion of the Universe is accelerating. Originally this was shown by studying Type Ia supernovae (SNe Ia) [1, 2, 3]. Measurements of the cosmic microwave background radiation (CMBR)[4] and baryon acoustic oscillations show the same thing [5].
The simplest and most common explanation is to assume that there is a cosmological constant $\Lambda$, such that the gravitational action normally proportional to the curvature scalar $R$ is modified to $R+2\Lambda$. The evidence [4] indicates that the universe is close to spatially flat, and the resulting $\Lambda$CDM model fits well with available data.
@@
To find $T$, it is first necessary to derive the stress-energy tensor from the matter Lagrangian $\mathcal{L}_{m}$. When dealing with ordinary matter or dark matter, the matter Lagrangian comes from either the standard model Lagrangian or whatever extension is responsible for dark matter, thus making the explicit form of the stress-energy tensor either complicated or unknown. A common strategy in both standard general relativity and in these modified theories is to assume a perfect fluid stress-energy tensor. The method for doing this in general relativity is well-known [8], but there are additional complications that occur in modified gravity, as pointed out by [9] and later elaborated by [10]. Multiple papers [7, 11, 12, 13, 14, 15], ignore these complications, making their conclusions suspect. One attempt to bypass the ambiguity is proposed by [16]. However, this attempt contained some sign inconsistencies, partially due to errors in the literature related to the metric conventions, and a detailed analysis of their work leads us to conclude that this approach is not productive.
@@
As pointed out by [10], in such theories, $f_{2}(T)$ can, in principle, be incorporated into the matter Lagrangian $\mathcal{L}_{m}$, and as such perhaps should not be considered as a modification of gravity at all. This is indeed the argument of [17]. However, whether this is considered as a modification of gravity or not, we can still ask the question of whether such a theory can account for the accelerating expansion of the Universe.
In this paper, we focus on the question of how one can reproduce a currently accelerating universe with a theory of the form Eq. (1). Since we are focusing on the effects of the stress-energy term, we use $f_{1}(R)=R$ and use the simplest possible form $f_{2}(T)=\lambda T^{\epsilon}$. As pointed out by [17], if we choose $\epsilon=2$, then standard model constraints at the weak scale places strong limits on $\lambda$. At the much lower energy densities relevant in current-day cosmology, the new term would be irrelevant compared to other terms in the action such as $\mathcal{L}_{m}$. This argument generalizes to conclude that any $\epsilon>1$ is irrelevant for present-day cosmology, and we will focus on $\epsilon<1$. The case when $\epsilon=0$ corresponds to the standard $\Lambda$CDM. For comparison we will choose $\epsilon=-1$, which should have an affect on the late-time Universe. We also assume the universe is flat, so as to not introduce unnecessary additional parameters. Thus our alternative has the same number of parameters as standard $\Lambda$CDM, since the cosmological constant $\Lambda$ is replaced by the new coupling constant $\lambda$.
In Section II, the $f(R,T)$ formalism is laid out, yielding the modified Einstein’s equations. In Section III, perfect fluids as discussed by [10, 8] are reviewed for the particular case of $f(R,T)=R+\lambda T^{\epsilon}$. Then in Section IV, the scale factor of the Universe as a function of time is derived for our model and compared to $\Lambda$CDM results. Finally in Section V, SNe Ia data from the Pantheon data set [18, 19, 20] is compared against both our model and $\Lambda$CDM.
@@
where the subscripts on $f$ denote partial derivatives [7]. Standard gravity with a cosmological constant $\Lambda$ can be recovered by choosing $f(R,T)=R+2\Lambda$. Note that the stress-energy is not conserved in these models [7]. Taking the divergence of Eq. 4 gives
@@ III Perfect Fluids
We will discover when considering non-standard gravity that the naïve number density and stress-energy tensor are not always conserved; nonetheless, we will treat Eq. (8) as a definition for the naïve pressure $p$.
In standard gravity, it is common to assume $\mathcal{L}_{m}=p$ [7], but it is worth understanding the origin of this expression, which for non-standard gravity turns out not to be so simple, as was pointed out by [8]. We start with the form
@@
$\qquad{}-\frac{1}{2\kappa^{2}}f^{\prime}_{2}(T)\frac{\partial\rho}{\partial s}(\rho+p)\;,$
@@
${\overline{T}}_{\mu\nu}$ $=T_{\mu\nu}^{\text{PF}}-g_{\mu\nu}\frac{\epsilon\lambda}{2\kappa^{2}{\overline{T}}^{1-\epsilon}+4\epsilon\lambda}n\frac{\partial}{\partial n}(\rho+p)\;,$ (16)
@@ IV The Scale Factor
The initial motivation for considering dark energy was the study of SNe Ia. A plot of the luminosity distance vs. red-shift $z$ seemed to indicate that, unlike a matter- or curvature-dominated universe, the universe was currently accelerating its expansion. It is useful to understand how this comes about in the standard $\Lambda$CDM model, with $f(R,T)=R+2\Lambda$. We assume a spatially flat universe, with metric
@@
where $a$ is an arbitrarily normalized scale factor. It is evident from Eq. (13a) that in this case the number density $n$ is conserved, and in the non-relativistic current universe we would have $\rho\propto n$, so that $\rho a^{3}$ will be a constant. The $tt$-component of Eq. (4) is then
@@
$a(t)=\begin{cases}N(3t)^{2/3}&t\rightarrow 0\\ NH_{\Lambda}^{-2/3}\exp(H_{\Lambda}t)&t\rightarrow\infty\;.\end{cases}$ (20)
Can we get a similar outcome from $f(R,T)$ gravity? We will assume again that the space metric is flat, given by Eq. (18). The situation is complicated because quantities like the energy and pressure derived just from the matter Lagrangian $\mathcal{L}_{m}$ will not be conserved, nor will be the number density. However, since the energy density $\rho$ and pressure $p$ are being derived simply from the matter Lagrangian, we expect that in the non-relativistic era, the pressure will still be $p=0$. We expect the naïve energy density $\rho\propto n$ because they are both derived from the standard matter Lagrangian $\mathcal{L}_{m}$.
@@
where the positive square root is favored so that ${\overline{T}}=\rho$ in the limit where the energy density is large or the coupling $\lambda$ is small. Due to how useful Eq. (21) is for rewriting the energy density in terms of the stress-energy trace, this equation will henceforth be used throughout our work without referencing it.
Assuming a flat FLRW metric, i.e. Eq. (18), the first Friedmann equation can be obtained in terms of ${\overline{T}}$ by combining Eqs. (4), (12), and (15) for $\epsilon=-1$ to yield
@@
We note that this is nearly identical to Eq. (20), the only difference being an overall scale increase in this model that is smaller by a factor of $(\sqrt{2}\alpha)^{-2/3}\approx 0.8021$. The behavior of the Hubble parameter is shown in Fig. 1. However, there is no particular reason to assume the asymptotic Hubble constants $H_{\Lambda}$ and $H_{\lambda}$ should match, since the goal of these models is not to make the Universe with particular future behavior, but instead to match the observed red-shift–luminosity curves.
@@ V Luminosity Distance and Observation
where $t_{0}$ is the time value for the present and $a_{0}=a(t_{0})$. In the $\Lambda$CDM-model,
@@
Figure 2: The distance modulus $\mu=m-M$ as a function of the red shift factor $z$. The points represent data taken from the Pantheon survey provide by [18]. The solid line corresponds to the best fit $\Lambda$CDM model. The dashed line represents the best fit of our model at $\tau_{0}=0.937$.
Fig. 2 shows the 1048 supernovae data from the Pantheon data set [18, 19, 20], superimposed with the best fit curves for both our model and $\Lambda$CDM. As is clear, the two theories for an appropriate choice of parameters yield virtually indistinguishable curves. To clarify what is going on, we have included Fig. 3, which shows the difference between $f(R,T)$ predictions and the best fit $\Lambda$CDM, for different values of $\lambda$. Also included in Fig. 3 are the binned supernovae data. Again, at least visually, it is not obvious which theory is better.
@@
As in the standard $\Lambda$CDM model, Hubble’s constant $H_{0}$ and the absolute magnitude $M$ for SNe Ia are degenerate parameters [18]. Recalling that for the $\epsilon=0$ case that the $f(R,T)$ model is equivalent to the $\Lambda$CDM model, the second parameter used in both theories is $\lambda$. In both cases there are 1046 degrees of freedom.
@@ VI Conclusion
We have studied the expansion of the Universe in $f(R,T)$ gravity with $f(R,T)=R+\lambda T^{-1}$ assuming a flat Universe. Like $\Lambda$CDM, we found an exponential growth transition from the matter dominated era to present day cosmology when considering the analysis of [10]. We also compared our predictions for luminosity distance versus redshift with SNe Ia data [18]. Our results show that this new model is competitive with, or even slightly better than, $\Lambda$CDM.
It is clear from our analysis that supernovae data indicate an accelerating Universe, but current data cannot necessarily distinguish the cause of such acceleration. Since our $f(R,T)$ model fits the data well, as evident in Fig. 3, the deviation between standard and non-standard cosmologies is most evident at high redshift so that more supernovae at high $z$ are needed to differentiate between these models.

### v2

@@ Abstract
The expansion of the Universe in $f(R,T)$ gravity is studied. By focusing on functions of the form $f(R,T)=f_{1}(R)+f_{2}(T)$, we assert that present-day acceleration can be achieved if the functional form of $f_{2}(T)$ either grows slowly or falls as a function of $T$. In particular, we demonstrate that when $f_{2}(T)\propto T^{-1}$, the Universe transitions to exponential growth at late times, just as it does in the standard cosmological model. A comparison of predictions of this model with type Ia supernovae shows that this model fits the data as well or even slightly better than the standard cosmological model without increasing the number of parameters.
@@ I Introduction
For over two decades, we have known that the expansion of the Universe is accelerating. Originally this was shown by studying type Ia supernovae (SNe Ia) [1, 2, 3]. Measurements of the cosmic microwave background radiation [4] and baryon acoustic oscillations show the same thing [5].
The simplest and most common explanation is to assume that there is a cosmological constant $\Lambda$, such that the gravitational action normally proportional to the curvature scalar $R$ is modified to $R+2\Lambda$. The evidence [4] indicates that the Universe is close to spatially flat, and the resulting $\Lambda$ cold dark matter ($\Lambda$CDM) model fits well with available data.
@@
To find $T$, it is first necessary to derive the stress-energy tensor from the matter Lagrangian $\mathcal{L}_{m}$. When dealing with ordinary matter or dark matter, the matter Lagrangian comes from either the standard model Lagrangian or whatever extension is responsible for dark matter, thus making the explicit form of the stress-energy tensor either complicated or unknown. A common strategy in both standard general relativity and in these modified theories is to assume a perfect fluid stress-energy tensor. The method for doing this in general relativity is well known [8], but there are additional complications that occur in modified gravity, as pointed out by [9] and later elaborated by [10]. Multiple papers [7, 11, 12, 13, 14, 15] ignore these complications, making their conclusions suspect. One attempt to bypass the ambiguity is proposed by [16]. However, this attempt contained some sign inconsistencies, partially due to errors in the literature related to the metric conventions, and a detailed analysis of their work leads us to conclude that this approach is not productive.
@@
As pointed out by [10], in such theories, $f_{2}(T)$ can, in principle, be incorporated into the matter Lagrangian $\mathcal{L}_{m}$ and, as such, perhaps should not be considered as a modification of gravity at all. This is indeed the argument of [17]. However, whether this is considered as a modification of gravity or not, we can still ask the question of whether such a theory can account for the accelerating expansion of the Universe.
In this paper, we focus on the question of how one can reproduce a currently accelerating Universe with a theory of the form of Eq. (1). Since we are focusing on the effects of the stress-energy term, we use $f_{1}(R)=R$ and use the simplest possible form $f_{2}(T)=\lambda T^{\epsilon}$. As pointed out by [17], if we choose $\epsilon=2$, then standard model constraints at the weak scale places strong limits on $\lambda$. At the much lower energy densities relevant in current-day cosmology, the new term would be irrelevant compared to other terms in the action such as $\mathcal{L}_{m}$. This argument generalizes to conclude that any $\epsilon>1$ is irrelevant for present-day cosmology, and we will focus on $\epsilon<1$. The case when $\epsilon=0$ corresponds to the standard $\Lambda$CDM. For comparison we will choose $\epsilon=-1$, which should have an effect on the late-time Universe. We also assume the Universe is flat, so as to not introduce unnecessary additional parameters. Thus our alternative has the same number of parameters as standard $\Lambda$CDM, since the cosmological constant $\Lambda$ is replaced by the new coupling constant $\lambda$.
In Sec. II, the $f(R,T)$ formalism is laid out, yielding the modified Einstein’s equations. In Sec. III, perfect fluids as discussed by [10, 8] are reviewed for the particular case of $f(R,T)=R+\lambda T^{\epsilon}$. Then in Sec. IV, the scale factor of the Universe as a function of time is derived for our model and compared to $\Lambda$CDM results. Finally in Sec. V, SNe Ia data from the Pantheon dataset [18, 19, 20] are compared against both our model and $\Lambda$CDM.
@@
where the subscripts on $f$ denote partial derivatives [7]. Standard gravity with a cosmological constant $\Lambda$ can be recovered by choosing $f(R,T)=R+2\Lambda$. Note that the stress-energy is not conserved in these models [7, 21] [^1] Note a missing term from [7] is corrected in [21]. Taking the divergence of Eq. (4) gives
@@ III Perfect Fluids
We will discover when considering nonstandard gravity that the naïve number density and stress-energy tensor are not always conserved; nonetheless, we will treat Eq. (8) as a definition for the naïve pressure $p$.
In standard gravity, it is common to assume $\mathcal{L}_{m}=p$ [7], but it is worth understanding the origin of this expression, which for nonstandard gravity turns out not to be so simple, as was pointed out by [8]. We start with the form
@@
$\qquad{}-\frac{1}{2\kappa^{2}}f^{\prime}_{2}(T)\frac{\partial}{\partial s}(\rho+p)\;,$
@@
${\overline{T}}_{\mu\nu}$ $=T_{\mu\nu}^{\text{PF}}-g_{\mu\nu}\frac{\epsilon\lambda}{2\kappa^{2}{\overline{T}}^{1-\epsilon}+4\epsilon\lambda}n\frac{\partial}{\partial n}(\rho+p)\;.$ (16)
@@ IV The Scale Factor
The initial motivation for considering dark energy was the study of SNe Ia. A plot of the luminosity distance versus redshift $z$ seemed to indicate that, unlike a matter- or curvature-dominated universe, the Universe was currently accelerating its expansion. It is useful to understand how this comes about in the standard $\Lambda$CDM model, with $f(R,T)=R+2\Lambda$. We assume a spatially flat Universe, with metric
@@
where $a$ is an arbitrarily normalized scale factor. It is evident from Eq. (13a) that in this case the number density $n$ is conserved, and in the nonrelativistic current Universe we would have $\rho\propto n$, so that $\rho a^{3}$ will be a constant. The $tt$-component of Eq. (4) is then
@@
$a(t)=\begin{cases}N(3t)^{2/3}&t\rightarrow 0\;,\\ NH_{\Lambda}^{-2/3}\exp(H_{\Lambda}t)&t\rightarrow\infty\;.\end{cases}$ (20)
Can we get a similar outcome from $f(R,T)$ gravity? We will assume again that the space metric is flat, given by Eq. (18). The situation is complicated because quantities like the energy and pressure derived just from the matter Lagrangian $\mathcal{L}_{m}$ will not be conserved, nor will be the number density. However, since the energy density $\rho$ and pressure $p$ are being derived simply from the matter Lagrangian, we expect that in the nonrelativistic era, the pressure will still be $p=0$. We expect the naïve energy density $\rho\propto n$ because they are both derived from the standard matter Lagrangian $\mathcal{L}_{m}$.
@@
where the positive square root is favored so that ${\overline{T}}=\rho$ in the limit where the energy density is large or the coupling $\lambda$ is small. Because of how useful Eq. (21) is for rewriting the energy density in terms of the stress-energy trace, this equation will henceforth be used throughout our work without referencing it.
Assuming a flat Friedmann-Lemaître-Robinson-Walker metric, i.e. Eq. (18), the first Friedmann equation can be obtained in terms of ${\overline{T}}$ by combining Eqs. (4), (12), and (15) for $\epsilon=-1$ to yield
@@
We note that this is nearly identical to Eq. (20), the only difference being an overall scale increase in this model that is smaller by a factor of $(\sqrt{2}\alpha)^{-2/3}\approx 0.8021$. The behavior of the Hubble parameter is shown in Fig. 1. However, there is no particular reason to assume the asymptotic Hubble constants $H_{\Lambda}$ and $H_{\lambda}$ should match, since the goal of these models is not to make the Universe with particular future behavior, but instead to match the observed redshift-luminosity curves.
@@ V Luminosity Distance and Observation
where $t_{0}$ is the time value for the present and $a_{0}=a(t_{0})$. In the $\Lambda$CDM model,
@@
Figure 2: The distance modulus $\mu=m-M$ as a function of the redshift factor $z$. The points represent data taken from the Pantheon survey provided by [18]. The solid line corresponds to the best fit $\Lambda$CDM model. The dashed line represents the best fit of our model at $\tau_{0}=0.937$.
Figure 2 shows the 1048 supernovae data from the Pantheon dataset [18, 19, 20], superimposed with the best fit curves for both our model and $\Lambda$CDM. As is clear, the two theories for an appropriate choice of parameters yield virtually indistinguishable curves. To clarify what is going on, we have included Fig. 3, which shows the difference between $f(R,T)$ predictions and the best fit $\Lambda$CDM, for different values of $\lambda$. Also included in Fig. 3 are the “binned” supernovae data. Again, at least visually, it is not obvious which theory is better.
@@
As in the standard $\Lambda$CDM model, Hubble’s constant $H_{0}$ and the absolute magnitude $M$ for SNe Ia are degenerate parameters [18]. Recalling that for the $\epsilon=0$ case the $f(R,T)$ model is equivalent to the $\Lambda$CDM model, the second parameter used in both theories is $\lambda$. In both cases there are 1046 degrees of freedom.
@@ VI Conclusion
We have studied the expansion of the Universe in $f(R,T)$ gravity with $f(R,T)=R+\lambda T^{-1}$ assuming a flat Universe. Like $\Lambda$CDM, we found an exponential growth transition from the matter-dominated era to present-day cosmology when considering the analysis of [10]. We also compared our predictions for luminosity distance versus redshift with SNe Ia data [18]. Our results show that this new model is competitive with, or even slightly better than, $\Lambda$CDM.
It is clear from our analysis that supernovae data indicate an accelerating Universe, but current data cannot necessarily distinguish the cause of such acceleration. Since our $f(R,T)$ model fits the data well, as evident in Fig. 3, the deviation between standard and nonstandard cosmologies is most evident at high redshift so that more supernovae at high $z$ are needed to differentiate between these models.
@@ Acknowledgments
Data Availability
The data that support the findings of this article are openly available [18].